# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v3

@@ Abstract
Adaptive Cruise Control (ACC) is a widely used driver assistance technology for maintaining desired speed and safe distance to the leading vehicle. This paper evaluates the security of the deep neural network (DNN) based ACC systems under runtime stealthy perception attacks that strategically inject perturbations into camera data to cause forward collisions. We present a context-aware strategy for the selection of the most critical times for triggering the attacks and a novel optimization-based method for the adaptive generation of image perturbations at runtime. We evaluate the effectiveness of the proposed attack using an actual vehicle, a publicly available driving dataset, and a realistic simulation platform with the control software from a production ACC system, a physical-world driving simulator, and interventions by the human driver and safety features such as Advanced Emergency Braking System (AEBS). Experimental results show that the proposed attack achieves 142.9 times higher success rate in causing hazards and 89.6% higher evasion rate than baselines, while being stealthy and robust to real-world factors and dynamic changes in the environment. This study highlights the role of human drivers and basic safety mechanisms in preventing attacks.
I Introduction
@@
Level-2 Advanced Driver Assistance Systems (ADAS) provide autonomous driving features while still requiring human attention at all times [18]. Examples include Adaptive Cruise Control (ACC) which controls longitudinal movement, Automatic Lane Centering (ALC) which controls lateral movement, and Advanced Emergency Braking System (AEBS) which controls braking through Automatic Emergency Braking (AEB) and provides warnings through Forward Collision Warning (FCW). Over 17 million passenger cars worldwide are equipped with ADAS [19].
@@
One important ADAS feature is ACC, which makes highway driving more comfortable by automatically changing the speed when traffic slows down or speeds up. ACC takes as input sensor measurements such as radar, Lidar, or camera and adjusts the speed to maintain a safe following distance to the lead vehicle [15, 12]. At the core of ACC lies the detection and tracking of the lead vehicle. Highly accurate methods [70, 80] for detection and tracking rely on Deep Learning (DL) based object detection using camera or fusion of camera and radar/Lidar data. A Longitudinal planner (LP) uses the prediction from the DL module to compute the desired speed and acceleration. Malfunctioning of the object detection module can have serious consequences including accidents. Given the critical role of object detection and tracking in the safety of ACC and that many commercial ACC systems (e.g., Tesla Autopilot [10], Comma.ai OpenPilot [29]) use DL-based object detection, these mechanisms must function correctly under any conditions, including in the presence of adversaries.
@@
Previous work has shown attacks against Deep Neural Networks (DNN) used for perception such as adversarial perturbations [95], adversarial patches [55, 51, 76] or well-crafted stickers on road signs [36], the road [84], or camera lenses [53] to change the predicted class or probability of detecting a target object or the lane lines. However, misprediction of the lane lines only affects the lateral control (ALC) and simply changing the lead object class or detection probability does not necessarily impact the LP enough to cause unsafe ACC behavior (e.g., sudden acceleration) [39]. Attacking DNN-based ACC systems necessitates influencing the relative distance and speed with respect to the lead vehicle. One such attack using physical adversarial patches to control the relative distance and speed to the lead vehicle was shown against a production ACC [39]. However, the attack required placing a conspicuous large patch on the back of a truck and driving the truck in front of the target vehicle, a method easily noticeable or preventable by human drivers (see Appendix D).
@@
Notably, previous work ignored the presence of safety interventions in the ADAS control loop. Although some attack works have considered AEB or FCW, they either do not apply to ACC [76] or directly change the state estimations [60]. Some recent works [97, 44] have used the system context to determine the optimum time for attack injection. However, they did not directly compute the perturbations for the DL perception module to affect relative speed and distance and could be easily detected by existing safety mechanisms or anomaly detection methods [96, 28].
To fill these gaps, in this paper, we focus on runtime stealthy attacks against DNN-based ACC systems that inject minimal image perturbations into the DNN input with the goal of causing the ACC controller to issue unsafe acceleration commands that cannot be mitigated by the human driver or existing safety mechanisms and lead to safety hazards, such as forward collision. The attacks are implemented as malware or malicious updates that are designed through offline analysis of the open-source code and data of a target ADAS and are integrated with the ADAS software through one-time access to a target vehicle or fleet of vehicles at the same time. We assume that the baseline ACC, ADAS software, and the existing safety mechanisms are trusted.
@@
  • We design a novel optimization-based approach and an adaptive algorithm to generate stealthy image perturbations and add them in the form of an adversarial patch to the input camera frames at runtime to fool DNN model and cause unsafe acceleration by ACC controller before being detected or mitigated by the human driver or the ADAS safety mechanisms.
  • We evaluate the effectiveness of the attacks with a real vehicle and driving dataset and a realistic simulation platform that integrates an open-source ADAS control software, OpenPilot from Comma.ai (with over 10,000 active users on the road) [29] and the state-of-the-art physical-world driving simulator, CARLA, with a driver reaction simulator and the typical ADAS safety mechanisms (AEB and FCW), which we implement based on the international standards.
  • Experimental results show that our context-aware attack strategy causes 28.6x more hazards than random attacks. The proposed optimization-based perturbation algorithm achieves a 100% attack success rate in four high-risk driving scenarios, 142.9x and 1.9x higher than random value perturbation and APGD-based methods [31]. Our approach is also stealthy and robust to real-world factors such as different camera positions, distances to lead vehicle, and weather and lighting conditions.
  • We observe similar results in the presence of human driver and safety feature interventions, where our attack achieves at least 18.7x higher success rate while being mitigated 89.6% less by the safety mechanisms compared to random and APGD-based attacks.
@@ II ADAS Overview
II-A Adaptive Cruise Control (ACC)
@@
The main goal of ACC is to maintain a safe following distance between the autonomous vehicle (referred to as Ego vehicle or AV) and the vehicle driving in the same lane in front of the Ego vehicle (referred to as lead vehicle or LV) by adjusting the Ego vehicle’s speed based on the estimated relative distance and relative speed to the lead vehicle.
Figure 1: ADAS architecture with ACC, AEBS, and ALC.
@@
Sensors. Existing DNN-based ACC systems either use camera data (e.g., Tesla Autopilot, Subaru Eyesight) or both camera and radar data (e.g., Apollo [21] and OpenPilot) to predict and track lead vehicles and objects. Other sensors such as GPS or IMU are also used to detect the current speed to match the target speed set by the human drivers.
@@
Lead Vehicle Detection. The most critical part of ACC is lead vehicle detection (LVD), which estimates the relative speed ($RS$) and distance ($RD$) to the lead vehicle using camera data or a fusion of camera and radar data. Sensor fusion is the process of combining measurements from multiple sensors (e.g., camera and radar) usually using a Kalman filter [22] to overcome the limitations of individual sensors and obtain a more accurate perception of the surrounding environment. Based on the LVD outputs, the main driving control actions (i.e., acceleration, deceleration, braking) are determined.
@@
Longitudinal Planner. The next stage involves determining the optimal speed based on the LVD outputs and current vehicle state. The longitudinal planner uses algorithms such as Model Predictive Control (MPC) to generate multiple desired speed trajectories, each representing a series of speeds in a certain following period [24].
@@
Vehicle Control. At each control cycle $t$, the plan from the longitudinal planner with the lowest speed and risk (e.g., risk of colliding with the lead vehicle) is selected by the ACC system and fed to a Propotional-Integral-Derivative (PID) [23] controller to get the specific optimal control command $u_{t}$ in the form of the throttle or brake amount such that the vehicle accurately and quickly follows the desired speed trajectory. Upon execution of the control command by the actuators, the vehicle’s physical state, $s_{t}$ (e.g., current speed, location), transitions to a new state $s_{t+1}$.
II-B ADAS Safety Mechanisms
The Advanced Emergency Braking System (AEBS), including FCW and AEB, is a fundamental safety mechanism that alerts drivers about potential collision risks with a lead obstacle and actively decelerates the vehicle to prevent accidents. Current ADAS predominantly incorporate AEBS as part of their safety mechanisms, albeit under different names. As shown in Fig. 1, most AEBS implementations utilize both camera and radar for collision prediction through sensor fusion [1] and make control actions based on the LVD outputs and other sensor measurements (see Appendix C). In addition, some safety principles (such as maximum acceleration limits as required by international standards, e.g., ISO 22179) or firmware safety checks (e.g., constraints on the output steering angle) are incorporated into the design of typical production ADAS to ensure driving safety [11].
Previous studies on security of autonmous driving either focus on Level 4 or fully autonomous vehicles without considering the impact of the human driver interventions during an emergency situation (e.g., abnormal acceleration) or have overlooked the inclusion of basic safety features like AEB or FCW and their impact on the attack effectiveness (see Table I). For a realistic assessment of ACC security, it is essential to evaluate the interventions of these safety features. At Level 2, drivers must maintain control and supervise ADAS functionalities [18]. There exists a research gap concerning how to make the combination of a human driver and an autonomous vehicle acceptably safe. The primary challenge lies in assessing the ability of human drivers to anticipate and respond to situations where automation may fail.
TABLE I: Attacks comparison for DNN perception or ADAS.
@@ II-B ADAS Safety Mechanisms
Method Type Vector Target AEBS Driver Level[18]
[36] Offline Stickers on road signs Classifier N N N/A
[58] Patch projected MTO N N L4
[84] Stickers on road ALC N N L2
[39] Patch on truck ACC N N L2
[76] Patch on road ALC Y N L2
[45] Runtime Perception inputs MTO N N L4
[44] Perception inputs MTO No FCW N L4
[60] Inner state variables FCW No AEB Y N/A
[97] Control commands ACC, ALC No AEB Y L2
@@
II-C OpenPilot
We use a production ADAS called OpenPilot from Comma.ai [29] as our case study. OpenPilot is the only open-source production Level-2 ADAS, designed with the goal of improving visual perception and automated control (with ACC and ALC) through installing an additional EON to the OBD-II port on a vehicle. The targeted ACC system in OpenPilot follows the typical DNN-based ACC system architecture described in Fig. 1 with an end-to-end system design [26]. Currently, OpenPilot supports over 250 car models (e.g., Toyota, Honda, etc.) [9], has more than 10,000 active users and has accumulated a total driving distance of over 100 million miles on actual roads [29]. It is reported to achieve state-of-the-art autonomous driving performance, beating 17 existing production ADAS on the market, including Tesla Autopilot and Audi Driver Assistance Plus in the overall ranking by Consumer Reports [30].
Comma.ai offers the community a closed-loop simulation environment [29] that integrates OpenPilot with a physical-world open urban driving simulator called CARLA, which can generate near-real high-quality camera image frames of the environment and has been widely used in the literature of autonomous driving [32]. However, in this default simulation, the sensor fusion only relies on the camera data since no radar sensor is available. Further, none of the typical ADAS safety mechanisms are included.
III Runtime Context-Aware Perception Attack
@@
III-A Attack Model
Attacker Objective. The objective of the attacker is to maximize the error in lead vehicle predictions by the LVD’s DNN module and cause forward collisions, while remaining stealthy to avoid being detected or prevented by the human driver or safety mechanisms (e.g., AEB, FCW).
The attack is crafted in a stealthy way such that it is not distinguishable from noise, human errors, or accidents. This enables it to remain hidden longer and makes it less easily detected by existing defense mechanisms. The attacker can accomplish this by targeting the DNN inputs ( 1 in Fig. 1) or directly manipulating the DNN outputs ( 2), depending on their capabilities and access level to the ACC system. In this paper, we primarily focus on DNN inputs to enhance stealthiness, as detailed in Section V-D.
Attacker Knowledge. We assume that the attacker gains comprehensive knowledge about the target ACC system design and implementation by reverse engineering a purchased or rented vehicle with identical control software as the victim vehicle [84, 46] or by studying publicly available documents or source code. This is possible given that some of the production ACC systems are open source [29, 21].

### v4

@@ Abstract.
Adaptive Cruise Control (ACC) is a widely used driver assistance technology for maintaining desired speed and safe distance to the leading vehicle. This paper evaluates the security of the deep neural network (DNN) based ACC systems under runtime stealthy perception attacks that strategically inject perturbations into camera data to cause forward collisions. We present a context-aware strategy for the selection of the most critical times for triggering the attacks and a novel optimization-based method for the adaptive generation of image perturbations at runtime. We evaluate the effectiveness of the proposed attack using an actual vehicle, a publicly available driving dataset, and a realistic simulation platform with the control software from a production ACC system, a physical-world driving simulator, and interventions by the human driver and safety features such as Advanced Emergency Braking System (AEBS). Experimental results show that the proposed attack achieves 142.9 times higher success rate in causing hazards and 82.6% higher evasion rate than baselines, while being stealthy and robust to real-world factors and dynamic changes in the environment. This study highlights the role of human drivers and basic safety mechanisms in preventing attacks.
Runtime Attack, Safety Intervention, AEBS, ADAS, ACC, DNN, Perception Attack, Stealthy.
journalyear: 2025copyright: rightsretainedconference: ACM Asia Conference on Computer and Communications Security; August 25–29, 2025; Hanoi, Vietnambooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’25), August 25–29, 2025, Hanoi, Vietnamdoi: 10.1145/3708821.3710832isbn: 979-8-4007-1410-8/25/08ccs: Computer systems organization Embedded and cyber-physical systemsccs: Security and privacy Systems security
1. Introduction
@@
Level-2 Advanced Driver Assistance Systems (ADAS) provide autonomous driving features while still requiring human attention at all times (SAEroadmap, ). Examples include Adaptive Cruise Control (ACC) which controls longitudinal movement, Automatic Lane Centering (ALC) which controls lateral movement, and Advanced Emergency Braking System (AEBS) which controls braking through Automatic Emergency Braking (AEB) and provides warnings through Forward Collision Warning (FCW). Over 17 million passenger cars worldwide are equipped with ADAS (canalys1, ).
@@
One important ADAS feature is ACC, which makes highway driving more comfortable by automatically changing the speed when traffic slows down or speeds up. ACC takes as input sensor measurements such as radar, Lidar, or camera and adjusts the speed to maintain a safe following distance to the lead vehicle (sayer2003adaptive, ; sae2018taxonomy, ). At the core of ACC lies the detection and tracking of the lead vehicle. Highly accurate methods (ravindran2020multi, ; sharma2018beyond, ) for detection and tracking rely on Deep Learning (DL) based object detection using camera or fusion of camera and radar/Lidar data. A Longitudinal planner (LP) uses the prediction from the DL module to compute the desired speed and acceleration. Malfunctioning of the object detection module can have serious consequences including accidents. Given the critical role of object detection and tracking in the safety of ACC and that many commercial ACC systems (e.g., Tesla Autopilot (autopilot, ), Comma.ai OpenPilot (commaai-openpilot, )) use DL-based object detection, these mechanisms must function correctly under any conditions, including in the presence of adversaries.
@@
Previous work has shown attacks against Deep Neural Networks (DNN) used for perception such as adversarial perturbations (zhou2022adversarial, ), adversarial patches (liu2018dpatch, ; lee2019physical, ; sato2021dirty, ) or well-crafted stickers on road signs (eykholt2018robust, ), the road (tencent2019experimental, ), or camera lenses (li2019adversarial, ) to change the predicted class or probability of detecting a target object or the lane lines. However, misprediction of the lane lines only affects the lateral control (ALC) and simply changing the lead object class or detection probability does not necessarily impact the LP enough to cause unsafe ACC behavior (e.g., sudden acceleration). Attacking DNN-based ACC systems necessitates influencing the relative distance and speed with respect to the lead vehicle. One such attack using physical adversarial patches to control the relative distance and speed to the lead vehicle was shown against a production ACC (guo2023adversarial, ). However, the attack required placing a conspicuous large patch on the back of a truck and driving the truck in front of the target vehicle, a method easily noticeable or preventable by human drivers (see Appendix D).
@@
Notably, previous work ignored the presence of safety interventions in the ADAS control loop. Although some attack works have considered AEB or FCW, they either do not apply to ACC (sato2021dirty, ) or directly change the state estimations (ma2021sequential, ). Some recent works (zhou2022strategic, ; jha2020ml, ) have used the system context to determine the optimum time for attack injection. However, they did not directly compute the perturbations for the DL perception module to affect relative speed and distance and could be easily detected by existing safety mechanisms or anomaly detection methods (zhou2023hybrid, ; Choi2020software, ).
To fill these gaps, in this paper, we focus on runtime stealthy attacks against DNN-based ACC systems that inject minimal image perturbations into the DNN input with the goal of causing ACC controller to issue unsafe acceleration commands that cannot be mitigated by the human driver or existing safety mechanisms and lead to safety hazards, such as forward collision. We assume baseline ACC, ADAS software, and existing safety mechanisms are trusted.
@@
  • We design a novel optimization-based approach and an adaptive algorithm to generate stealthy image perturbations and add them in the form of an adversarial patch to the input camera frames at runtime to fool DNN model and cause unsafe acceleration by ACC controller before being detected or mitigated by the human driver or ADAS safety mechanisms.
  • We evaluate the effectiveness of the attacks with a real vehicle and driving dataset and a realistic simulation platform that integrates an open-source ADAS control software, OpenPilot from Comma.ai (with over 10,000 active users on the road) (commaai-openpilot, ) and the state-of-the-art physical-world driving simulator, CARLA, with a driver reaction simulator and the typical ADAS safety mechanisms (AEB and FCW), which we implement based on the international standards.
  • Experimental results show that our context-aware attack strategy causes 28.6x more hazards than random attacks. The proposed optimization-based perturbation algorithm achieves a 100% attack success rate in four high-risk driving scenarios (in the absence of safety interventions), 142.9x and 1.9x higher than random value perturbation and APGD-based methods (croce2020reliable, ). Our approach is also stealthy and robust to real-world factors such as different camera positions, distances to lead vehicle, and weather and lighting conditions.
  • We observe similar results in the presence of the human driver and safety feature interventions, where our attack still achieves an 82.6% success rate, while all the random and APGD-based attacks are mitigated by the safety mechanisms.
@@ 2. ADAS Overview
2.1. Adaptive Cruise Control (ACC)
@@
The main goal of ACC is to maintain a safe following distance between the autonomous vehicle (referred to as Ego vehicle or AV) and the vehicle driving in the same lane in front of the AV (referred to as lead vehicle or LV) by adjusting the AV speed based on the estimated relative distance and relative speed to the LV.
Figure 1. ADAS architecture with ACC, AEBS, and ALC.
@@
Sensors. Existing DNN-based ACC systems either use camera data (e.g., Tesla Autopilot, Subaru Eyesight) or both camera and radar data (e.g., Apollo (Apollo, ) and OpenPilot) to predict and track LVs and objects. Other sensors, such as GPS or IMU, are also used to detect current speed to match the target speed set by the drivers.
@@
Lead Vehicle Detection. The most critical part of ACC is lead vehicle detection (LVD), which estimates the relative speed ($RS$) and distance ($RD$) to the LV using camera data or a fusion of camera and radar data. Sensor fusion is the process of combining measurements from multiple sensors (e.g., camera and radar) usually using a Kalman filter (bishop2001kalman, ) to overcome the limitations of individual sensors and obtain a more accurate perception of the surrounding environment. Based on the LVD outputs, the main driving control actions (i.e., acceleration, deceleration, braking) are determined.
@@
Longitudinal Planner. The next stage involves determining the optimal speed based on the LVD outputs and current vehicle state. The longitudinal planner uses algorithms such as Model Predictive Control (MPC) to generate multiple desired speed trajectories, each representing a series of speeds in a certain following period (camacho2013model, ).
@@
Vehicle Control. At each control cycle $t$, the plan from the longitudinal planner with the lowest speed and risk (e.g., risk of colliding with the LV) is selected by the ACC system and fed to a Propotional-Integral-Derivative (PID) (bishop2011modern, ) controller to get the specific optimal control command $u_{t}$ in the form of the throttle or brake amount such that the vehicle accurately and quickly follows the desired speed trajectory. Upon execution of the control command by the actuators, the vehicle’s physical state, $s_{t}$ (e.g., current speed, location), transitions to a new state $s_{t+1}$.
Table 1. Comparison of attacks on DNN-based ADAS.
@@
Method Type Vector Target AEBS Driver Level(SAEroadmap, )
(eykholt2018robust, ) Offline Stickers on road signs Classifier N N N/A
(ma_wip_2023, ) Patch projected MTO N N L4
(tencent2019experimental, ) Stickers on road ALC N N L2
(guo2023adversarial, ) Patch on truck ACC N N L2
(sato2021dirty, ) Patch on road ALC Y N L2
(jia2019fooling, ) Runtime Perception inputs MTO N N L4
(jha2020ml, ) Perception inputs MTO No FCW N L4
(ma2021sequential, ) Inner state variables FCW No AEB Y N/A
(zhou2022strategic, ) Control commands ACC, ALC No AEB Y L2
@@
2.2. ADAS Safety Mechanisms
The Advanced Emergency Braking System (AEBS), including FCW and AEB, is a fundamental ADAS safety mechanism that alerts drivers about potential collision risks with a lead obstacle and actively decelerates the vehicle to prevent accidents. As shown in Fig. 1, most AEBS implementations utilize both camera and radar for collision prediction through sensor fusion (AdasSensorFusion, ) and make control actions based on the LVD outputs and other sensor measurements (see Appendix C). In addition, some safety principles (such as maximum acceleration limits as required by international standards, e.g., ISO 22179) or firmware safety checks (e.g., constraints on the output steering angle) are incorporated into the design of typical production ADAS to ensure driving safety (openpilot_safety, ).
Previous studies on security of autonmous driving either focus on Level 4 or fully autonomous vehicles without considering the impact of the human driver interventions during an emergency situation (e.g., abnormal acceleration) or have overlooked the inclusion of basic safety features like AEB or FCW and their impact on the attack effectiveness (see Table 1). For a realistic assessment of ACC security, it is essential to evaluate the interventions of these safety features. At Level 2, drivers must maintain control and supervise ADAS functionalities (SAEroadmap, ). There exists a research gap concerning how to make the combination of a human driver and an autonomous vehicle acceptably safe. The primary challenge lies in assessing the ability of human drivers to anticipate and respond to situations where automation may fail.
2.3. OpenPilot
We use a production ADAS called OpenPilot from Comma.ai (commaai-openpilot, ) as our case study. OpenPilot is the only open-source production Level-2 ADAS, designed with the goal of improving visual perception and automated control (with ACC and ALC) through installing custom hardware to the OBD-II port on a vehicle. The targeted ACC system in OpenPilot follows the typical DNN-based ACC system architecture described in Fig. 1 with an end-to-end system design (chen2022level, ). Currently, OpenPilot supports over 250 car models (e.g., Toyota, Honda, etc.) (Carmodels-OpenPilot, ), has more than 10,000 active users and has accumulated a total driving distance of over 100 million miles on actual roads (commaai-openpilot, ). It is reported to achieve state-of-the-art autonomous driving performance, beating 17 existing production ADAS on the market in overall ranking by Consumer Reports (CR-OpenPilot, ).
The DNN model used by OpenPilot, called Supercombo, utilizes an EfficientNet-B2 based CNN model to process image data (schmedding2024strategic, ). It incorporates the state of the vehicle and the environment by adding additional inputs from traffic conventions and the desired state. Multiple branches of GEMM (General Matrix Multiply) operations are then used to derive various predictions, such as lane lines, LVs, and vehicle pose, resulting in a total of 6,472 outputs (Supercombo, ).
Comma.ai offers the community a closed-loop simulation environment (commaai-openpilot, ) that integrates OpenPilot with a physical-world open urban driving simulator called CARLA, which can generate near-real high-quality camera image frames of the environment and has been widely used in the literature of autonomous driving (Dosovitskiy17, ). However, in this default simulation, the sensor fusion only relies on the camera data since no radar sensor is available. Further, none of the typical ADAS safety mechanisms are included.
3. Runtime Context-Aware Perception Attack
@@
3.1. Attack Model
Attacker Objective. The objective of the attacker is to maximize the error in LV predictions by the LVD’s DNN module and cause forward collisions, while remaining stealthy to avoid being detected or prevented by driver or safety mechanisms (e.g., AEBS).
The attack is crafted in a stealthy way such that it is not distinguishable from noise, human errors, or accidents. This enables it to remain hidden longer and makes it less easily detected/prevented by existing defense mechanisms (e.g., anomaly detection (AVattacksurvey2021, ) or input transformation (zhang2019defending, ; dziugaite2016study, ; xu2017feature, ; xu2017feature, )). The attacker can accomplish this by targeting the DNN inputs ( 1 in Fig. 1) or directly manipulating the DNN outputs ( 2), depending on their capabilities and access level to the ACC system. In this paper, we primarily focus on DNN inputs to enhance stealthiness, as detailed in Section 5.4.
Attacker Knowledge. We assume the attacker gains comprehensive knowledge about the target ACC system design and implementation by reverse engineering a purchased or rented vehicle with identical control software as the victim vehicle (tencent2019experimental, ; AVattacksurvey2021, ) or by studying publicly available documents or source code. This is possible given some production ACC systems are open source (commaai-openpilot, ; Apollo, ).